OpenStudy (anonymous):

how can i put this into y= form -x+4y=-20

OpenStudy (mathteacher1729):

step 0 write the original problem -x+4y=-20 step 1 add x to both sides 4y = x - 20 step 2 divide by four on both sides y = (1/4)x - 5 step 3 party. You are done. :p

OpenStudy (anonymous):

so the final answer is y=(1/4)x-5?

OpenStudy (mathteacher1729):

Yup. Did the steps make sense? Do you feel you can do this for similar problems now?

OpenStudy (anonymous):

yes thank you so much. that was SUPER helpful. :) much like having a teacher right infront of me :)

OpenStudy (mathteacher1729):

Yaay! :D